OpenStudy (anonymous):

how do i find the length of the transverse and conjugate axis of (x-4)^2/36 - (y-2)^2/9=1?

OpenStudy (anonymous):

@ranga

OpenStudy (amistre64):

consider when x=4, and also when y=2 only one of them those produces a viable solution ... the under parts tells the length of half the axis

OpenStudy (anonymous):

is there a certain equation i would use to find the length of the transverse and conjugate axis?

OpenStudy (amistre64):

(x-4)^2/36 - (y-2)^2/9=1 ^^^ when this is zero - (y-2)^2/9=1 has no solution (x-4)^2/36 - (y-2)^2/9=1 ^^^ when this is zero (x-4)^2/36 =1 has 2 solutions ... we want to observe that 36

OpenStudy (amistre64):

half the transverse axis is equal to sqrt(36) half the conjugate would be the sqrt(9) by default

OpenStudy (amistre64):

errr, im almost confusing that with an ellipse :/

OpenStudy (anonymous):

uh oh :o

OpenStudy (amistre64):